### v2

# On the number of generalized cospectral mates of graphs

Abstract
This paper establishes an upper bound on the number of generalized cospectral mates of simple graphs, where the generalized spectrum consists of the spectrum of a graph and its complement. Moving beyond the classical problem of identifying graphs determined by their generalized spectrum, we address the more quantitative question of how many non-isomorphic graphs can share the same generalized spectrum. Our approach is based on arithmetic constraints derived from the Smith Normal Form (SNF) of the walk matrix, which leads to a tight upper bound on the number of generalized cospectral mates of a graph. Our upper bound applies to a much broader class of graphs than those previously shown to have no generalized cospectral mates (graphs determined by generalized spectrum). Consequently, this work extends the family of graphs for which strong and informative spectral uniqueness results are available.
Keywords: Adjacency matrix, Walk matrix, Cospectral graphs, Generalized spectrum
1 Introduction
Let $G=(V,E)$ be a simple graph with $n$ vertices, meaning it is finite, undirected, and contains no loops or multiple edges. The adjacency matrix $A(G)$ is the $n\times n$ symmetric matrix where each entry $a_{ij}=1$ if vertices $v_{i}$ and $v_{j}$ are adjacent, and $a_{ij}=0$ otherwise. The spectrum of a graph $G$, denoted $\operatorname{Spec}(G)$, is the multiset of eigenvalues of its adjacency matrix $A(G)$, counted with algebraic multiplicity.
Two graphs $G$ and $H$ are isomorphic, denoted $G\cong H$, if there exists a bijection $f:V(G)\to V(H)$ such that vertices $u$ and $v$ are adjacent in $G$ if and only if $f(u)$ and $f(v)$ are adjacent in $H$. In other words, isomorphic graphs have identical structure, differing only by a relabeling of vertices. The complement of a graph $G$, denoted $\overline{G}$, is the graph on the same vertex set where two vertices are adjacent if and only if they are not adjacent in $G$. The adjacency matrix of the complement is given by $A(\overline{G})=J-I-A(G)$, where $J$ is the $n\times n$ all-ones matrix and $I$ is the identity matrix.
A graph $G$ is said to be determined by its spectrum (DS) if every graph with the same spectrum as $G$ is isomorphic to $G$. A central open question in spectral graph theory is: “Which graphs are DS?” This question was first raised by Günthard and Primas [2] in the context of chemistry and is closely related to Kac’s famous question [4], “Can one hear the shape of a drum?” Fisher [1] modeled the drum as a graph whose sound is described by its spectrum, linking Kac’s question to the graph being DS. Haemers conjectured that almost all graphs are DS [9, 3]. Nevertheless, proving that a given graph is DS remains challenging; see [9, 10] for surveys.
Wang and Xu [11] considered this problem in the context of generalized spectrum, defined as the pair $(\operatorname{Spec}(G),\operatorname{Spec}(\overline{G}))$. Two or more graphs are generalized cospectral if they have the same generalized spectrum. While isomorphic graphs are clearly generalized cospectral, the converse does not hold in general. Pairwise non-isomorphic graphs that are generalized cospectral are referred to as generalized cospectral mates. A graph is said to be determined by its generalized spectrum (DGS) if it has no generalized cospectral mate.
Recent advances have leveraged the properties of the walk matrix to address these questions. For a graph $G$ with $n$ vertices, the walk matrix $W(G)$ is the $n\times n$ matrixa with columns $e,\,A(G)e,\,\dots,\,A(G)^{n-1}e$, where $e$ is the all-ones column vector of length $n$. A graph $G$ is controllable if $W(G)$ is non-singular. Wang [12] introduced a large family of controllable graphs and proved that all graphs in this family are DGS. Extending these ideas to oriented graphs, Qiu et al. [6] introduced a family of oriented graphs determined by their generalized skew spectra.
A natural extension of this line of research addresses a fundamental question: how many generalized cospectral mates can a graph have? Wang et al. [13] identified a family of graphs where each graph has at most one generalized cospectral mate and proposed an algorithm to find such a mate. Later on, they provided an algorithm to find generalized cospectral mates for controllable or almost controllable graphs [14]. However, the practical feasibility of this algorithm is constrained by the computational difficulty of factorizing large integers. In the context of oriented graphs, Raza et al. [8] and Lin et al. [5] independently established upper bounds on the number of their generalized cospectral mates for a large family.
In this paper, we introduce a large family of graphs and establish an upper bound on the number of their generalized cospectral mates. Let $d_{1}(M)\mid d_{2}(M)\mid\dots\mid d_{n}(M)$ denote the invariant factors of the Smith Normal Form of a matrix $M$. We define $\mathcal{F}_{n}$ as the family of controllable graphs on $n$ vertices where each graph $G$ satisfies:
$d_{\lceil n/2\rceil}(W(G))=1\quad\text{and}\quad d_{n-1}(W(G))=2.$ (1)
To analyze the relationship between generalized cospectral mates in $\mathcal{F}_{n}$, we utilize the concept of the level of a rational matrix:
Definition 1.
The level of a rational matrix $Q$, denoted $\ell(Q)$, is defined as the smallest positive integer $x$ such that $xQ$ is an integral matrix.
For graphs in $\mathcal{F}_{n}$, we prove the following theorem:
Theorem 1.
Let $G\in\mathcal{F}_{n}$, and let $H_{1}$ and $H_{2}$ be graphs that are generalized cospectral with $G$. Let $Q_{1}$ and $Q_{2}$ be the rational orthogonal matrices satisfying $Q_{i}^{\mathrm{T}}A(G)Q_{i}=A(H_{i})$ and $Q_{i}e=e$ for $i=1,2$. If $\ell(Q_{1})=\ell(Q_{2})$, then $H_{1}\cong H_{2}$.
We will later establish the existence of the orthogonal matrices referenced in Theorem 1. This theorem shows that, for graphs in $\mathcal{F}_{n}$, each possible level of a rational orthogonal matrix yields at most one generalized cospectral mate. Consequently, the number of generalized cospectral mates is bounded above by the number of possible levels:
Theorem 2.
Let $G\in\mathcal{F}_{n}$. Let $d_{n}(W(G))=\prod_{j=1}^{m}p_{j}^{k_{j}}$ be the prime factorization of the last invariant factor of the walk matrix $W(G)$. Then $G$ can have at most $\left(\prod_{j=1}^{m}k_{j}\right)-1$ generalized cospectral mates.
We begin by reviewing essential preliminaries in Section 2. In Section 3, we investigate the properties of a specific set of orthogonal matrices; these properties establish the framework for proving our main results in Section 4. We then provide illustrative examples and numerical analysis in Section 5 to demonstrate the application of our results. Section 6 concludes the paper with a summary of our findings and potential directions for future research.
2 Preliminaries
In this section, we review the essential results required for the proofs of Theorem 1 and 2. A matrix $Q$ is called regular if $Qe=e$ where $e$ is the all-ones column vector. The following theorem characterizes graphs sharing the same generalized spectrum:
Theorem 3.
[11] Let $G$ be a controllable graph and $H$ be a graph generalized cospectral with $G$. Then $H$ must be controllable and there exists a unique regular rational orthogonal matrix $Q$ such that:
$Q^{\mathrm{T}}A(G)Q=A(H),$ (2)
Moreover, the unique matrix $Q$ satisfies $Q=W(G)W(H)^{-1}$.
Motivated by this theorem, we consider the set of orthogonal matrices that generate a valid adjacency matrix. Let $O_{n}(\mathbb{Q})$ denote the set of regular orthogonal matrices of order $n$ with rational entries. We define:
$\Gamma(G)=\{Q\in O_{n}(\mathbb{Q})\mid Q^{\mathrm{T}}A(G)Q=A(H)\text{ for graph }H\},$
The level of a matrix $Q\in\Gamma(G)$ provides a crucial link between spectral properties and graph isomorphism. Specifically, if $\ell(Q)=1$, then $Q$ is an integral orthogonal matrix. Since $Q$ is also regular, it must be a permutation matrix. Consequently, if Eq. (2) holds with a matrix of level one, the graphs $G$ and $H$ are isomorphic. Therefore, generalized cospectral mates arise only when $\ell(Q)>1$. The next result characterizes the relationship between rational orthogonal matrices that generate isomorphic graphs.
Lemma 4.
[8, 13] Let $G$ be a graph, and let $H_{1}$ and $H_{2}$ be graphs generalized cospectral with $G$. Let $Q_{1},Q_{2}\in\Gamma(G)$ be the corresponding rational orthogonal matrices such that $Q_{i}^{\mathrm{T}}A(G)Q_{i}=A(H_{i})$ and $Q_{i}e=e$ for $i=1,2$. If $Q_{2}=Q_{1}P$, where $P$ is a permutation matrix, then $H_{1}\cong H_{2}$.
In this paper, our strategy is to demonstrate that if two matrices in $\Gamma(G)$ have the same level, they must differ by only by column permutations. The following lemma provides a simple criterion for bounding the level of a rational matrix based on modular congruences.
Lemma 5.
[8] Let $Q$ be rational matrix, and let $x$ and $k$ be positive integers. If $xQ$ is an integral matrix and $xQ\equiv 0\pmod{k}$, then $\ell(Q)\mid\frac{x}{k}$.
An important tool in this context is the Smith Normal Form (SNF) of integral matrices. An integral matrix $V$ of order $n$ is called unimodular if $\det{V}=\pm 1$. For a prime $p$, we denote the finite field of order $p$ by $\mathbb{F}_{p}$, and the rank of an integral matrix $M$ over this field by $\operatorname{rank}_{p}(M)$.
Theorem 6.
For a full-rank integral matrix $M$, there exist unimodular matrices $V_{1}$ and $V_{2}$ such that $M=V_{1}SV_{2}$, where $S=\operatorname{diag}(d_{1}(M),d_{2}(M),\ldots,d_{n}(M))$ is the SNF of $M$, and the invariant factors satisfy $d_{i}(M)\mid d_{i+1}(M)$ for $i=1,2,\ldots,n-1$.
The determinant of an integral $M$ can thus be written as
$\det{M}=\pm\prod_{i=1}^{n}d_{i}(M).$ (3)
For a prime $p$, $\operatorname{rank}_{p}(M)$ is equal to the number of invariant factors of $M$ not divisible by $p$. Consequently, the definition of $\mathcal{F}_{n}$ implies that for any $G\in\mathcal{F}_{n}$, we have $\operatorname{rank}_{p}(W(G))=n-1$ for every odd prime $p$ dividing $d_{n}(W(G))$, and $\operatorname{rank}_{2}(W(G))\geq\lceil\frac{n}{2}\rceil$. The following lemma provides a general upper bound for the rank of the walk matrix over $\mathbb{F}_{2}$.
Lemma 7.
[12] Let $G$ be a graph with $n$ vertices. Then $\operatorname{rank}_{2}(W(G))\leq\left\lceil\frac{n}{2}\right\rceil$.
Combining Lemma 7 with the lower bound derived above, we conclude that $\operatorname{rank}_{2}(W(G))=\left\lceil\frac{n}{2}\right\rceil$ for every $G\in\mathcal{F}_{n}$. The next lemma establishes a connection between the level of a rational orthogonal matrix and the last invariant factors of the integral matrices it relates.
Lemma 8.
[7] Let $X$ and $Y$ be two non-singular integral matrices such that $QX=Y$, where $Q$ is a rational orthogonal matrix. Then $\ell(Q)$ divides the greatest common divisor of $d_{n}(X)$ and $d_{n}(Y)$, where $d_{n}(X)$ and $d_{n}(Y)$ are the $n$-th invariant factors of $X$ and $Y$, respectively.
For any matrix $Q\in\Gamma(G)$, we have $Q^{\mathrm{T}}W(G)=W(H)$ for some graph $H$ by Theorem 3. It follows that $\ell(Q)$ must divide $\operatorname{gcd}(d_{n}(W(G)),d_{n}(W(H)))$, and therefore it must divide $d_{n}(W(G))$. However, certain arithmetic properties of $W(G)$ can restrict which divisors of $d_{n}(W(G))$ are actually attainable as levels for matrices in $\Gamma(G)$.
For integers $m$, $k$, and $n$ (with $n\neq 0$), we write $m^{k}\parallel n$ to indicate that $m^{k}$ divides $n$ exactly; that is, $m^{k}\mid n$ but $m^{k+1}\nmid n$. The unique integer $k$ satisfying this condition is denoted by $\operatorname{ord}_{m}(n)$. The following lemma establishes that the highest power of an odd prime dividing $d_{n}(W(G))$ cannot divide $\ell(Q)$.
Lemma 9.
[7] Let $G$ be a graph, let $Q\in\Gamma(G)$ and let $p$ be an odd prime. If $p^{k}\parallel\det{W(G)}$ and $\operatorname{rank}_{p}(W(G))=n-1$, then $p^{k}\nmid\ell(Q)$ or equivalently, $\ell(Q)=\frac{d_{n}(W(G))}{p}$.
3 Arithmetic Constraints on Matrices in $O_{n}(\mathbb{Q})$
In this section, we establish several properties of the matrices in $O_{n}(\mathbb{Q})$ that are necessary for the proof of Theorem 1. We begin by identifying conditions under which the solution space of a linear system modulo $p^{k}$ is one-dimensional.
Lemma 10.
Let $M$ be a full-rank integral matrix of order $n$. If a prime $p$ and integer $k\geq 1$ satisfy $p^{k}\mid d_{n}(M)$ and $\operatorname{rank}_{p}(M)=n-1$, then every solution to $Mv\equiv 0\pmod{p^{k}}$ is a scalar multiple of a fixed vector $w\not\equiv 0\pmod{p}$ over the ring $\mathbb{Z}/p^{k}\mathbb{Z}$.
Proof.
Let $M=V_{1}SV_{2}$ be the SNF of $M$. Since $V_{1}$ is unimodular, it is invertible over $\mathbb{Z}/p^{k}\mathbb{Z}$. Therefore, the solutions to $Mv\equiv 0\pmod{p^{k}}$ are equivalent to the solutions of
$SV_{2}v\equiv 0\pmod{p^{k}}$.
Let $v^{\prime}=V_{2}v$. The system becomes $Sv^{\prime}\equiv 0\pmod{p^{k}}$.
Since $\operatorname{rank}_{p}(M)=n-1$, it follows that $p\nmid d_{i}(M)$ for $i=1,2,\ldots,n-1$. Thus, the first n1n-1 entries of $v^{\prime}$ must be congruent to zero modulo $p^{k}$. The last entry of $v^{\prime}$, however, can take any value in $\mathbb{Z}/p^{k}\mathbb{Z}$. Therefore, the solutions to $Sv^{\prime}\equiv 0\pmod{p^{k}}$ are of the form
$v^{\prime}\equiv yw^{\prime}\pmod{p^{k}}$.
where $y$ is any integer in $\mathbb{Z}/p^{k}\mathbb{Z}$ and $w^{\prime}=[0,\dots,0,1]^{\mathrm{T}}$.
Since $V_{2}$ is unimodular, the inverse $V_{2}^{-1}$ exists and is also unimodular. Multiplying both sides of the relation $v^{\prime}=V_{2}v$ by $V_{2}^{-1}$, we obtain $v=V_{2}^{-1}v^{\prime}$. Substituting the solution $v^{\prime}\equiv yw^{\prime}\pmod{p^{k}}$, we obtain:
$v\equiv yV_{2}^{-1}w^{\prime}\pmod{p^{k}}$.
Define $w=V_{2}^{-1}w^{\prime}$. Note that $w$ is the last column of the matrix $V_{2}^{-1}$. We claim that $w\not\equiv 0\pmod{p}$. If $w\equiv 0\pmod{p}$, then every entry in the last column of $V_{2}^{-1}$ would be divisible by $p$. Consequently, the determinant of $V_{2}^{-1}$ would be divisible by $p$. This contradicts the fact that $\det(V_{2}^{-1})=\pm 1$. Hence, $w\not\equiv 0\pmod{p}$, and every solution $v$ to $Mv\equiv 0\pmod{p^{k}}$ is a scalar multiple of $w$ over $\mathbb{Z}/p^{k}\mathbb{Z}$. ∎
The following lemmas establish that if vectors satisfying specific constraints are scalar multiples over the ring $\mathbb{Z}/p^{k}\mathbb{Z}$, then their inner product must be congruent to zero modulo $p^{2k}$.
Lemma 11.
Let $u$ and $v$ be $n$-dimensional integer vectors, and let $p$ be an odd prime and $k$ be a positive integer. If the following conditions hold:
  1. 1.
    $u\equiv\alpha v\pmod{p^{k}}$ where $\alpha$ is an integer not divisible by $p$,
  2. 2.
    $u^{\mathrm{T}}u\equiv v^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$,
then $u^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$.
Proof.
From condition (1), we have:
$u-\alpha v\equiv 0\pmod{p^{k}}$,
Taking the inner product of both sides with itself gives
$(u-\alpha v)^{\mathrm{T}}(u-\alpha v)\equiv 0\pmod{p^{2k}}$,
which expands to
$u^{\mathrm{T}}u-2\alpha u^{\mathrm{T}}v+\alpha^{2}v^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$.
Since $u^{\mathrm{T}}u\equiv 0\pmod{p^{2k}}$ and $v^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$, we have
$2\alpha u^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$.
As $p$ is odd and $\alpha$ is not divisible by $p$, it follows that $2\alpha$ is invertible in $\mathbb{Z}/p^{2k}\mathbb{Z}$. Hence $u^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$.
Lemma 12.
Let $u$ and $v$ be $n$-dimensional integer vectors, and let $k$ be a positive integer. If the following conditions hold:
  1. 1.
    $u\equiv\alpha v\pmod{2^{k}}$ where $\alpha$ is an odd integer,
  2. 2.
    $\operatorname{ord}_{2}(u^{\mathrm{T}}u)=\operatorname{ord}_{2}(v^{\mathrm{T}}v)=2k$,
  3. 3.
    $\operatorname{ord}_{2}(u^{\mathrm{T}}e)=\operatorname{ord}_{2}(v^{\mathrm{T}}e)=k$,
then $u^{\mathrm{T}}v\equiv 0\pmod{2^{2k}}$.
Proof.
Based upon our assumptions, we have: (i) $u^{\mathrm{T}}u=2^{2k}a$, (ii) $v^{\mathrm{T}}v=2^{2k}b$. (iii) $u^{\mathrm{T}}e=2^{k}c$, (iv) $v^{\mathrm{T}}e=2^{k}d$, where $a$, $b$, $c$, and $d$ are odd integers. From condition (1), we have
$u-\alpha v=2^{k}m,$ (4)
for some integer vector $m$. Multiplying each side by $e^{\mathrm{T}}$ gives
$e^{\mathrm{T}}u-\alpha e^{\mathrm{T}}v=2^{k}e^{\mathrm{T}}m.$
Substituting the values of $e^{\mathrm{T}}u$ and $e^{\mathrm{T}}v$ gives
$2^{k}c-\alpha(2^{k}d)=2^{k}e^{\mathrm{T}}m,$
which simplifies to
$c-\alpha d=e^{\mathrm{T}}m.$
Note that since $c$ and $\alpha d$ are both odd, their difference $c-\alpha d$ is even. Thus, $e^{\mathrm{T}}m$ is even. Moreover, we can express $m^{\mathrm{T}}m$ as
$m^{\mathrm{T}}m$ $=\sum_{i=1}^{n}m_{i}^{2}$
$\equiv\sum_{i=1}^{n}m_{i}$
$\equiv e^{\mathrm{T}}m\pmod{2}$,
where $m_{i}$ is the $i^{\text{th}}$ entry of $m$. Since $e^{\mathrm{T}}m$ is even, it follows that $m^{\mathrm{T}}m$ is also even.
Now, taking the inner product of both sides of Eq. (4) with itself gives:
$(u-\alpha v)^{\mathrm{T}}(u-\alpha v)$ $=(2^{k}m)^{\mathrm{T}}(2^{k}m)$
$u^{\mathrm{T}}u-2\alpha u^{\mathrm{T}}v+\alpha^{2}v^{\mathrm{T}}v$ $=2^{2k}m^{\mathrm{T}}m.$
Substituting the values of $u^{\mathrm{T}}u$ and $v^{\mathrm{T}}v$ gives:
$2^{2k}a-2\alpha u^{\mathrm{T}}v+\alpha^{2}(2^{2k}b)$ $=2^{2k}m^{\mathrm{T}}m$
$-2\alpha u^{\mathrm{T}}v$ $=2^{2k}(m^{\mathrm{T}}m-(a+\alpha^{2}b))$
$-\alpha u^{\mathrm{T}}v$ $=2^{2k-1}(m^{\mathrm{T}}m-(a+\alpha^{2}b)).$
Note that since $a$ and $\alpha^{2}b$ are both odd, their sum $a+\alpha^{2}b$ is even. Since $m^{\mathrm{T}}m$ is also even, it follows that $m^{\mathrm{T}}m-(a+\alpha^{2}b)$ is even. Thus, we conclude
$\alpha u^{\mathrm{T}}v\equiv 0\pmod{2^{2k}}$.
Since $\alpha$ is odd, it is invertible in $\mathbb{Z}/2^{2k}\mathbb{Z}$. Thus, $u^{\mathrm{T}}v\equiv 0\pmod{2^{2k}}$. ∎
In the next lemma, we explore the properties of the column vectors of matrices in $O_{n}(\mathbb{Q})$
Lemma 13.
Let $Q\in O_{n}(\mathbb{Q})$, $\bar{Q}=\ell(Q)Q$ and $v_{1},v_{2},\dots,v_{n}$ be the column vectors of $\bar{Q}$. Let $p$ be a prime factor of $\ell(Q)$ and let $k=\operatorname{ord}_{p}(\ell(Q))$. Then:
  1. 1.
    $v_{i}^{\mathrm{T}}v_{j}\equiv 0\pmod{p^{2k}}$ for all $i,j$;
  2. 2.
    $p^{2k}\parallel v_{i}^{\mathrm{T}}v_{i}$ for all $i$;
  3. 3.
    $p^{k}\parallel v_{i}^{\mathrm{T}}e$ for all $i$.
Proof.
Let $\ell=\ell(Q)$. Since $Q$ is orthogonal, $Q^{\mathrm{T}}Q=I$, and therefore $\bar{Q}^{\mathrm{T}}\bar{Q}=(\ell Q)^{\mathrm{T}}(\ell Q)=\ell^{2}I.$ As the $(i,j)$-th entry of $\bar{Q}^{\mathrm{T}}\bar{Q}$ is $v_{i}^{\mathrm{T}}v_{j}$, it follows that
$v_{i}^{\mathrm{T}}v_{j}=\begin{cases}\ell^{2}&\text{if }i=j,\\ 0&\text{if }i\neq j.\end{cases}$
Since $k=\operatorname{ord}_{p}(\ell)$, we have $p^{k}\parallel\ell$ and thus $p^{2k}\parallel\ell^{2}$. It follows that $v_{i}^{\mathrm{T}}v_{j}\equiv 0\pmod{p^{2k}}$ for all $i,j$ while $p^{2k}\parallel v_{i}^{\mathrm{T}}v_{i}$ for all $i$. This proves (1) and (2).
To prove (3), since $Q$ is regular and orthogonal, we have $Q^{\mathrm{T}}e=e$. Multiplying by $\ell$ gives $\bar{Q}^{\mathrm{T}}e=\ell e.$ The $i^{\text{th}}$ component of $\bar{Q}^{\mathrm{T}}e$ is $v_{i}^{\mathrm{T}}e=\ell$. Since $p^{k}\parallel\ell$, we conclude that $p^{k}\parallel v_{i}^{\mathrm{T}}e$. ∎
We now prove that if the columns of two matrices in $O_{n}(\mathbb{Q})$ are scalar multiples of a common vector over the ring $\mathbb{Z}/p^{k}\mathbb{Z}$, then $p$ does not divide the level of their product.
Lemma 14.
Let $Q_{1},Q_{2}\in O_{n}(\mathbb{Q})$ and let $p$ be a prime. Let $\bar{Q}_{1}=\ell(Q_{1})Q_{1}$ and $\bar{Q}_{2}=\ell(Q_{2})Q_{2}$. If the following conditions hold:
  1. 1.
    $\operatorname{ord}_{p}(\ell(Q_{1}))=\operatorname{ord}_{p}(\ell(Q_{2}))=k$, for some integer $k\geq 1$,
  2. 2.
    The column vectors of both $\bar{Q}_{1}$ and $\bar{Q}_{2}$ are scalar multiples of a common vector $w\not\equiv 0\pmod{p}$ over $\mathbb{Z}/p^{k}\mathbb{Z}$,
then $p\nmid\ell(Q_{1}^{\mathrm{T}}Q_{2})$.
Proof.
Let $R=\bar{Q}_{1}^{\mathrm{T}}\bar{Q}_{2}$ be an integral matrix with entries $r_{ij}=u_{i}^{\mathrm{T}}v_{j}$, where $u_{i}$ and $v_{j}$ denote the $i^{\text{th}}$ and $j^{\text{th}}$ columns of $\bar{Q}_{1}$ and $\bar{Q}_{2}$, respectively. We show that $r_{ij}\equiv 0\pmod{p^{2k}}$ for all $i,j$.
By the definition of level, there exist column vectors $u^{\prime}$ of $\bar{Q}_{1}$ and $v^{\prime}$ of $\bar{Q}_{2}$ such that $u^{\prime}\not\equiv 0\pmod{p}$ and $v^{\prime}\not\equiv 0\pmod{p}$. Since both $u^{\prime}$ and $v^{\prime}$ are also scalar multiples of $w$ over $\mathbb{Z}/p^{k}\mathbb{Z}$, there exists an integer $\alpha\not\equiv 0\pmod{p}$ such that:
$u^{\prime}\equiv\alpha v^{\prime}\pmod{p^{k}}$, (5)
Furthermore, since every column vector $u_{i}$ and $v_{j}$ is a scalar multiple of $w$ over $\mathbb{Z}/p^{k}\mathbb{Z}$, we can express them as:
$u_{i}\equiv c_{i}u^{\prime}+p^{k}x_{i}\pmod{p^{2k}}$, (6)
$v_{j}\equiv d_{j}v^{\prime}+p^{k}y_{j}\pmod{p^{2k}}$, (7)
where $c_{i},d_{j}$ are integers and $x_{i},y_{j}$ are integer vectors.
Since $Q_{1}$ and $Q_{2}$ satisfy the conditions of Lemma 13, it follows that $u_{i}^{\mathrm{T}}u^{\prime}\equiv 0\pmod{p^{2k}}$ and $v_{j}^{\mathrm{T}}v^{\prime}\equiv 0\pmod{p^{2k}}$ for each column $u_{i}$ and $v_{j}$. Moreover, $\operatorname{ord}_{p}({u^{\prime}}^{\mathrm{T}}u^{\prime})=\operatorname{ord}_{p}({v^{\prime}}^{\mathrm{T}}v^{\prime})=2k$ and $\operatorname{ord}_{p}({u^{\prime}}^{\mathrm{T}}e)=\operatorname{ord}_{p}({v^{\prime}}^{\mathrm{T}}e)=k$. Therefore, we can apply Lemma 11 (for odd primes) or Lemma 12 (for $p=2$) to establish that ${u^{\prime}}^{\mathrm{T}}v^{\prime}\equiv 0\pmod{p^{2k}}$.
Multiplying Eq. (6) by ${u^{\prime}}^{\mathrm{T}}$ yields:
${u^{\prime}}^{\mathrm{T}}u_{i}$ $\equiv{u^{\prime}}^{\mathrm{T}}(c_{i}u^{\prime}+p^{k}x_{i})\pmod{p^{2k}}$
$\equiv c_{i}{u^{\prime}}^{\mathrm{T}}u^{\prime}+p^{k}{u^{\prime}}^{\mathrm{T}}x_{i}\pmod{p^{2k}}$
As ${u^{\prime}}^{\mathrm{T}}u_{i}\equiv 0\pmod{p^{2k}}$ and $c_{i}{u^{\prime}}^{\mathrm{T}}u^{\prime}\equiv 0\pmod{p^{2k}}$, it follows that $p^{k}{u^{\prime}}^{\mathrm{T}}x_{i}\equiv 0\pmod{p^{2k}}$, which implies ${u^{\prime}}^{\mathrm{T}}x_{i}\equiv 0\pmod{p^{k}}$. Using Eq. (5), we get ${u^{\prime}}^{\mathrm{T}}x_{i}\equiv\alpha{v^{\prime}}^{\mathrm{T}}x_{i}\pmod{p^{k}}$. As $\alpha\not\equiv 0\pmod{p}$, we conclude that ${v^{\prime}}^{\mathrm{T}}x_{i}\equiv 0\pmod{p^{k}}$. Using a similar approach, we can prove that ${u^{\prime}}^{\mathrm{T}}y_{j}\equiv 0\pmod{p^{k}}$ for each $y_{j}$.
Finally, for any columns $u_{i}$ and $v_{j}$, we have:
$u_{i}^{\mathrm{T}}v_{j}$ $\equiv(c_{i}u^{\prime}+p^{k}x_{i})^{\mathrm{T}}(d_{j}v^{\prime}+p^{k}y_{j})$
$\equiv c_{i}d_{j}{u^{\prime}}^{\mathrm{T}}v^{\prime}+c_{i}p^{k}{u^{\prime}}^{\mathrm{T}}y_{j}+d_{j}p^{k}x_{i}^{\mathrm{T}}v^{\prime}+p^{2k}x_{i}^{\mathrm{T}}y_{j}\pmod{p^{2k}}$.
We established that ${u^{\prime}}^{\mathrm{T}}y_{j}\equiv 0\pmod{p^{k}}$ and ${v^{\prime}}^{\mathrm{T}}x_{i}\equiv 0\pmod{p^{k}}$. Since ${u^{\prime}}^{\mathrm{T}}v^{\prime}\equiv 0\pmod{p^{2k}}$, every term on the right-hand side vanishes modulo $p^{2k}$. Thus, we conclude that $u_{i}^{\mathrm{T}}v_{j}\equiv 0\pmod{p^{2k}}$ for all $i,j$, which implies $R\equiv 0\pmod{p^{2k}}$. Substituting the definition of $R$, we have
$\ell(Q_{1})\ell(Q_{2})Q_{1}^{\mathrm{T}}Q_{2}\equiv 0\pmod{p^{2k}}$.
As $Q_{1}^{\mathrm{T}}Q_{2}$ is a rational matrix, by Lemma 5, it follows that $\ell(Q_{1}^{\mathrm{T}}Q_{2})\mid\frac{\ell(Q_{1})\ell(Q_{2})}{p^{2k}}$. Since $p^{k}\parallel\ell(Q_{1})$ and $p^{k}\parallel\ell(Q_{2})$, it follows that $p^{2k}\parallel\ell(Q_{1})\ell(Q_{2})$. Consequently, $p\nmid\frac{\ell(Q_{1})\ell(Q_{2})}{p^{2k}}$, which implies $p\nmid\ell(Q_{1}^{\mathrm{T}}Q_{2})$, completing the proof. ∎
Building up on this, the following lemma establishes that if matrices in $O_{n}(\mathbb{Q})$ having the same level satisfy certain conditions, then their product must be a permutation matrix.
Lemma 15.
Let $Q_{1},Q_{2}\in O_{n}(\mathbb{Q})$ such that $\ell(Q_{1})=\ell(Q_{2})=\ell$, and let $\bar{Q}_{1}=\ell Q_{1}$ and let $\bar{Q}_{2}=\ell Q_{2}$. Let
$\ell=\prod_{j=1}^{m}p_{j}^{k_{j}}$
be the prime factorization of $\ell$. Suppose that for each $j\in\{1,\dots,m\}$, there exists a vector $w_{j}\not\equiv 0\pmod{p_{j}}$ such that columns of both $\bar{Q}_{1}$ and $\bar{Q}_{2}$ is a scalar multiple of $w_{j}$ over the ring $\mathbb{Z}/p_{j}^{k_{j}}\mathbb{Z}$. Then $Q_{1}^{\mathrm{T}}Q_{2}$ is a permutation matrix.
Proof.
Since $\bar{Q}_{1}$ and $\bar{Q}_{2}$ are integral, the product $\bar{Q}_{1}^{\mathrm{T}}\bar{Q}_{2}=\ell^{2}Q_{1}^{\mathrm{T}}Q_{2}$ is an integral matrix. It follows from the definition of level that $\ell(Q_{1}^{\mathrm{T}}Q_{2})$ divides $\ell^{2}$.For each prime factor $p_{j}$ of $\ell$, the assumptions satisfy the conditions of Lemma 14 with $k=k_{j}$. Thus, $p_{j}\nmid\ell(Q_{1}^{\mathrm{T}}Q_{2})$. Since $\ell(Q_{1}^{\mathrm{T}}Q_{2})$ divides $\ell^{2}$ but contains no prime factor of $\ell$, it must be that $\ell(Q_{1}^{\mathrm{T}}Q_{2})=1$. Therefore, $Q_{1}^{\mathrm{T}}Q_{2}$ is an integral orthogonal matrix.
Finally, since $Q_{1},Q_{2}\in O_{n}(\mathbb{Q})$, it follows that $Q_{1}^{\mathrm{T}}Q_{2}\in O_{n}(\mathbb{Q})$. An integral orthogonal matrix with row sums equal to 1 must be non-negative, which implies $Q_{1}^{\mathrm{T}}Q_{2}$ is a permutation matrix. ∎
4 Proofs of Theorem 1 and Theorem 2
We begin with defining some auxiliary matrices that are critical for the proof of Theorem 1:
Lemma 16.
[12] Let $\phi(x)=x^{n}+c_{1}x^{n-1}+\cdots+c_{n-1}x+c_{n}$ be the characteristic polynomial of a graph $G$ with adjacency matrix $A=A(G)$. Define
$M(G):=\begin{cases}A^{n/2}+c_{2}A^{n/2-1}+\cdots+c_{n-2}A+c_{n}I,&\text{if $n$ is even},\\ A^{(n+1)/2}+c_{2}A^{(n-1)/2}+\cdots+c_{n-3}A^{2}+c_{n-1}A,&\text{if $n$ is odd}.\end{cases}$
Then $M(G)e\equiv 0\pmod{2}$, where $e$ is the all-ones vector.
Using this result, the integral matrix $\hat{W}(G)$ is defined as follows:
Definition 2.
Let $G$ be a graph with the adjacency matrix $A=A(G)$ and the matrix $M=M(G)$. $\hat{W}(G)$ is defined as:
$\hat{W}(G):=\left[\,e,\,Ae,\,\ldots,\,A^{\lceil n/2\rceil}e,\,\frac{Me}{2},\,\frac{AMe}{2},\,\ldots,\,\frac{A^{n-\lceil n/2\rceil-1}Me}{2}\,\right].$
As observed by Qiu et al. [7], if $H$ is generalized cospectral to $G$ and $Q\in\Gamma(G)$ such that $Q^{\mathrm{T}}A(G)Q=A(H)$, then $Q^{\mathrm{T}}M(G)e=M(H)e$ and consequently, $Q^{\mathrm{T}}\hat{W}(G)=\hat{W}(H)$. The following lemma relates the SNF of $\hat{W}(G)$ to that of the original walk matrix $W(G)$.
Lemma 17.
[7] Let $G$ be a graph with walk matrix $W(G)$. If $\operatorname{rank}_{2}{W(G)}=\lceil\frac{n}{2}\rceil$, then for each $1\leq i\leq n$, the invariant factors of $\hat{W}(G)$ satisfy:
$d_{i}(\hat{W}(G))=\begin{cases}d_{i}(W(G))&\text{if }1\leq i\leq\left\lceil\frac{n}{2}\right\rceil,\\ \frac{d_{i}(W(G))}{2}&\text{if }\left\lceil\frac{n}{2}\right\rceil<i\leq n.\end{cases}$
The next lemma establishes the $\operatorname{rank}_{p}(\hat{W}(G))$ for every prime factor $p$ of $d_{n}(\hat{W}(G))$ when $G\in\mathcal{F}_{n}$.
Lemma 18.
Let $G\in\mathcal{F}_{n}$, then for every prime factor $p$ of $d_{n}(\hat{W}(G))$, we have $\operatorname{rank}_{p}(\hat{W}(G))=n-1$.
Proof.
As $G\in\mathcal{F}_{n}$, we established earlier that $\operatorname{rank}_{2}(W(G))=\lceil\frac{n}{2}\rceil$. Therefore, by Lemma 17, the invariant factors of $\hat{W}(G)$ are given by:
$d_{i}(\hat{W}(G))=\begin{cases}1&\text{if }1\leq i\leq n-1,\\ \frac{d_{n}(W(G))}{2}&\text{if }i=n.\end{cases}$
It follows that $\operatorname{rank}_{p}(\hat{W}(G))=n-1$ for every prime $p$ dividing $d_{n}(\hat{W}(G))$. This completes the proof. ∎
We can now proceed to the proof of Theorem 1.
Proof of Theorem 1.
Let $\bar{Q}_{i}=\ell Q_{i}$ for $i=1,2$ be the associated integral matrices. Let the prime factorization of $\ell$ be given by
$\ell=\prod_{j=1}^{m}p_{j}^{k_{j}}.$
Since $Q_{i}^{\mathrm{T}}\hat{W}(G)=\hat{W}(H_{i})$, transposing the equation gives $\hat{W}(G)^{\mathrm{T}}Q_{i}=\hat{W}(H_{i})^{\mathrm{T}}$. Multiplying by $\ell$, we obtain $\hat{W}(G)^{\mathrm{T}}\bar{Q}_{i}=\ell\hat{W}(H_{i})^{\mathrm{T}}$. As $\hat{W}(H)$ is an integral matrix and $\ell\equiv 0\pmod{p_{j}^{k_{j}}}$ for each prime $p_{j}$ dividing $\ell$, it follows that $\hat{W}(G)^{\mathrm{T}}\bar{Q}_{i}\equiv 0\pmod{p_{j}^{k_{j}}}$. By Lemma 18, we have $\operatorname{rank}_{p_{j}}(\hat{W}(G)^{\mathrm{T}})=\operatorname{rank}_{p_{j}}(\hat{W}(G))=n-1$.
Consequently, for each $p_{j}$, the columns of $\bar{Q}_{i}$ are the solutions to a linear system satisfying the conditions of Lemma 10. Hence, there exists a vector $w_{j}\not\equiv 0\pmod{p_{j}}$ such that the columns of both $\bar{Q}_{1}$ and $\bar{Q}_{2}$ are scalar multiples of $w_{j}$ over the ring $\mathbb{Z}/p_{j}^{k_{j}}\mathbb{Z}$.
The conditions of Lemma 15 are therefore satisfied, implying that $Q_{1}^{\mathrm{T}}Q_{2}=P$ for some permutation matrix $P$. It follows that $Q_{2}=Q_{1}P$. By Lemma 4, we conclude that $H_{1}$ is isomorphic to $H_{2}$. ∎
The proof of Theorem 2 follows as a direct consequence of Theorem 1 by counting the possible levels of matrices in $\Gamma(G)$.
Proof of Theorem 2.
As 2 must be a factor of $d_{n}(W(G))$, let the prime factorization of $d_{n}(W(G))$ be written as $2^{k_{1}}\prod_{j=2}^{m}p_{j}^{k_{j}}$, where $p_{1}=2$ and $p_{2},\dots,p_{m}$ are odd primes. Consider a matrix $Q\in\Gamma(G)$. Since $Q^{\mathrm{T}}\hat{W}(G)=\hat{W}(H)$ for some graph $H$ generalized cospectral to $G$, Lemma 8 implies that $\ell(Q)\mid d_{n}(\hat{W}(G))$. By Lemma 17, we have $d_{n}(\hat{W}(G))=\frac{d_{n}(W(G))}{2}$. Consequently, the multiplicity of the factor 2 in $d_{n}(\hat{W}(G))$ is $k_{1}-1$, which implies that $2^{k_{1}}\nmid\ell(Q)$.
Moreover, for each odd prime $p_{j}$ ($j\geq 2$), by Lemma 9, we know that $p_{j}^{k_{j}}\nmid\ell(Q)$. It follows that for each prime $p_{j}$, the exponent satisfies $\operatorname{ord}_{p_{j}}(\ell(Q))\in\{0,1,\dots,k_{j}-1\}$. Thus, there are exactly $k_{j}$ possible choices for the exponent of each prime factor. Let $L(G)$ denote the set of all possible values for $\ell(Q)$. The total number of these values is the product of the number of choices for each prime factor:
$|L(G)|=\prod_{j=1}^{m}k_{j}.$
We must exclude the trivial case of the level $\ell=1$, which corresponds to graphs isomorphic to $G$. By Theorem 1, each valid level corresponds to at most one generalized cospectral mate of $G$. Therefore, the number of generalized cospectral mates of $G$ is at most
$|L(G)|-1=\left(\prod_{j=1}^{m}k_{j}\right)-1$
This completes the proof. ∎
5 An Example and Numerical Evaluation
Figure 1: Graph $G$ and all of its generalized cospectral mates $H_{1},H_{2}$ and $H_{3}$.
In this section, we present an example of a graph in $\mathcal{F}_{n}$ alongside experimental results.
Example 1.
Let $n=10$. Consider the graph $G$ shown in Fig. 1, with the adjacency matrix:
$A(G)=\begin{pmatrix}0&0&0&0&0&1&0&1&0&0\\ 0&0&0&0&0&0&1&0&1&1\\ 0&0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&0&1&1&1\\ 0&0&0&0&0&0&0&1&1&0\\ 1&0&0&0&0&0&0&0&1&1\\ 0&1&1&0&0&0&0&1&0&1\\ 1&0&0&1&1&0&1&0&0&1\\ 0&1&0&1&1&1&0&0&0&1\\ 0&1&0&1&0&1&1&1&1&0\end{pmatrix}$
The invariant factors of the walk matrix $W(G)$.
$S=\operatorname{diag}(1,1,1,1,1,2,2,2,2,19350).$
As these factors satisfy Eq. (1), it follows that $G\in\mathcal{F}_{n}$. The prime factorization of the last invariant factor $d_{n}(W(G))$ is:
$d_{n}(W(G))=2\times 3\times 3\times 5\times 5\times 43$
According to Theorem 2, $G$ can have at most $3$ generalized cospectral mates. By a brute-force search over all graphs of size 1010, we find that $G$ has exactly 33 generalized cospectral mates, denoted $H_{1},H_{2},$ and $H_{3}$ (see Fig. 1). Thus, $G$ is a tight example attaining the bound given in Theorem 2.
Let $Q_{1},Q_{2},Q_{3}\in\Gamma(G)$ be the matrices such that:
$A(H_{1})=Q_{1}^{T}A(G)Q_{1},\quad A(H_{2})=Q_{2}^{T}A(G)Q_{2},\quad A(H_{3})=Q_{3}^{T}A(G)Q_{3}.$
These matrices are computed using Theorem 3. We find that their distinct levels are $\ell(Q_{1})=3$, $\ell(Q_{2})=15$, and $\ell(Q_{3})=5$, which satisfy the conditions of Theorem 1.
$n$ Controllable $|\mathcal{F}_{n}|$
11 6,705 3,573
12 7,708 3,646
13 8,441 3,844
14 9,061 3,895
15 9,427 3,911
20 9,982 3,979
25 9,998 3,897
30 10,000 3,973
35 10,000 3,930
40 10,000 3,953
45 10,000 3,908
50 10,000 3,920
Table 1: Statistics for 10,000 random graphs ($p=0.5$) for each order $n$. The columns report, respectively, the order $n$, the number of controllable graphs and the number graphs in $\mathcal{F}_{n}$.
In addition, we conducted a numerical experiment involving 10,000 random graphs generated with an edge probability of $0.5$ for various values of $n$. For each $n$, we counted the number of graphs belonging to $\mathcal{F}_{n}$. For smaller values of $n$, we also computed the bound given in Theorem 2, denoted by $\alpha(G)$. The results, summarized in Table 1, indicate that as $n$ increases, approximately $39\%$ of the generated graphs belong to $\mathcal{F}_{n}$.
6 Conclusion
In this paper, we have addressed the problem of determining the number of generalized cospectral mates for a graph by analyzing the arithmetic properties of the walk matrix. We identified a significant family of controllable graphs, denoted by $\mathcal{F}_{n}$, and established a concrete upper bound on the number of their generalized cospectral mates.
By leveraging the Smith Normal Form of the walk matrix and the level of rational orthogonal matrices, we proved that distinct levels yield non-isomorphic mates, allowing us to bound the number of mates based on the prime factorization of the last invariant factor of $W(G)$. Our numerical experiments demonstrate that $\mathcal{F}_{n}$ encompasses a substantial proportion of random graphs, approximately $39\%$ as $n$ increases, and we provided a tight example where the derived bound is attained.
Several avenues for future research emerge from this work. First, the definition of $\mathcal{F}_{n}$ relies on specific constraints on the invariant factors, namely that $d_{\lceil n/2\rceil}=1$ and $d_{n-1}=2$. A natural progression would be to relax these conditions to cover a broader range of controllable graphs. Generalizing the modular constraints on the rational orthogonal matrices could extend the applicability of these bounds to families where the rank of the walk matrix over finite fields behaves differently.
Furthermore, the algebraic techniques developed here need not be limited to the adjacency matrix. Future investigations could explore establishing similar upper bounds in other settings, such as using the signless Laplacian or the distance matrix. Finally, while this study focused on the walk matrix generated by the all-ones vector $e$, employing different control input vectors could reveal new structural insights. This approach might facilitate the determination of generalized cospectrality for graphs that are not controllable with respect to $e$ but may be controllable with respect to other vectors.